OpenStudy (anonymous):

the absolute global maximum and absolute function value of x^3 - 3x^2 + 1 on the interval (-.5,4)

OpenStudy (australopithecus):

what are you asking?

OpenStudy (australopithecus):

please clarify

OpenStudy (anonymous):

well im given these options a) 225,49/16 b)2, 0 c)17,-3 d)1,-3 e) 17, 1/8

OpenStudy (anonymous):

i need to find the absolute min and max values of the function on that interval

OpenStudy (australopithecus):

first take the derivative of the function

OpenStudy (anonymous):

3(x-2)x

OpenStudy (australopithecus):

f(x) = x^3 - 3x^2 + 1 f'(x) = 3x^(2) - 6x

OpenStudy (australopithecus):

so we are looking on the interval (-.5,4) Step 1 - First check the domain of the derivative what is it?

OpenStudy (australopithecus):

The domain is obviously \[\mathbb{R}\] so we dont have any critical points there Step 2 - Set f'(x) = 0 what values of x make the equation = 0

OpenStudy (anonymous):

its b

OpenStudy (anonymous):

2 and zero

OpenStudy (anonymous):

ok annonymous, can you just give me a quick rundown how

OpenStudy (australopithecus):

0 = 3(x-2)x we notice x = 0 x = 2 make the equation equal to 0 Are both numbers in the interval: (-.5,4) Yes they are so we will use both numbers to define the function in the interval So remember f'(x) > 0 the function is increasing f'(x) < 0 the function is decreasing so test numbers that are in the intervals of (-0.5, 0), (0 2), and (2, 4) to see if the function is positive or negative. Best to make a chart 

OpenStudy (anonymous):

first find x when f'(x)=0 f'(x)=3x^2-6x 3x^2-6x=0 3x(x-2)=0 so we'll have critical points at x=0,2, that also means that we have absolute extremes at those points which are exactly the absolute max and min.

OpenStudy (anonymous):

ok that was nice thanks

OpenStudy (australopithecus):

so lets test interval (-0.5, 0) f'(-0.2) =3((-0.2)-2)(-0.2) so 3(-2.2)(-0.2) = a negative number thus f'(x) < 0 thus the function is decreasing